OpenStudy (anonymous):

the perpindicular to 2x-4y=8 passin thourgh the point (1<-2. wat is the slope intercpt form

OpenStudy (aroub):

u mean the points (1,-2) right??

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

sry

OpenStudy (aroub):

its okay!! :)

OpenStudy (aroub):

-4y=-2x+8 y=1/2-2 y=mx+b , m=-2 , (1,-2) -2=-2(1)+b -2=-2+b -2+2=b 0=b y=-2x

OpenStudy (aroub):

oh sry slope intercept form ---> y+2=-2(x-1)